OpenStudy (anonymous):

If a snowball melts so that its surface area decreases at a rate of 10 cm2/min, find the rate at which the diameter decreases when the diameter is 12 cm.

zepdrix (zepdrix):

The Surface Area of a sphere is,\[\large A=4\pi r^2\] And they gave us the rate at which the Surface Area is decreasing,\[\large A'=-10\] We want to write our Surface Area formula in terms of Diameter. It is currently written in terms of the Radius. How can we change that? :)

OpenStudy (anonymous):

r = d/2 putting it into the equation gives \[A = 4\pi d ^{2} \div 4\] \[A = \pi d ^{2} \] Take derivative in both side \[A \prime = 2 \pi d d \prime\] d' = -10/ 2 pi*12